OpenStudy (anonymous):

What is the axis of symmetry for the graph of y = 4x2 + 16x + 10? x = 4 x = −4 x = −2 x = 2 Question 2 (Multiple Choice Worth 1 points) [9.03] What is the domain of the graph of y = −x2 + 6x + 11? x ≤ 3 All real numbers x ≤ −3 x ≥ 3

OpenStudy (anonymous):

You are looking for the center line of the parabola... x=-b/(2a)=-16/(2x4).

OpenStudy (anonymous):

can u help me :(

OpenStudy (anonymous):

And all quadratic functions have a domain from negative infinity to positive infinity.

OpenStudy (anonymous):

so for the first answer i got -2 im i right

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

yeahhhh thanks WAT about the 2 question

OpenStudy (anonymous):

Look up to my second reply.

OpenStudy (anonymous):

x ≤ −3

OpenStudy (anonymous):

did i get it right

OpenStudy (anonymous):

I'm trying to use some math lingo so that I am not just spitting out the answer. It is not x<=-3

OpenStudy (anonymous):

ohh its x ≤ 3

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

from negative infinity to positive infinity

OpenStudy (anonymous):

x ≥ 3 or all real numbers i think its x ≥ 3 thats what i got

OpenStudy (anonymous):

Domain asks the question, which numbers that I put in will give me an answer that exists? (or all possible x values) There is nothing that will cause a problem with this (no divide by zeros or negative square roots) so any thing will work.

OpenStudy (anonymous):

ok thanks so its all real numbers

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Which equation below has an axis of symmetry of x = 4 over 3? y = −3x2 − 4x + 5 y = −3x2 + 4x + 5 y = −3x2 − 8x − 1 y = −3x2 + 8x − 1 Question 6 (Multiple Choice Worth 1 points) [9.03] What is the range of the graph of y = −3x2 − 12x + 5? All real numbers y ≤ −2 y ≥ 17 y ≤ 17

OpenStudy (anonymous):

For the first one. You need to check -b/(2a) for each until you find 4/3 as your answer.

OpenStudy (anonymous):

For the second one: Range is all of the possible y values. It takes a bit more analysis than domain. First find where the vertex will occur (x value) using -b/(2a). I get -2. Then put that number back into the function as x to find the y... that is where the parabola starts (I get -3(-2)(-2)-12(-2)+5=17). Then see if it is right side up or upside down. This one is opening downward because of the minus sign in front... so for y-values (range), it starts at 17 and goes down.

OpenStudy (anonymous):

thanks